# A Benchmark for End-to-End Zero-Shot Biomedical Relation Extraction with LLMs: Experiments with OpenAI Models

### v1

@@ Abstract
Objective: Zero-shot methodology promises to cut down on costs of dataset annotation and domain expertise needed to make use of NLP. Generative large language models trained to align with human goals have achieved high zero-shot performance across a wide variety of tasks. As of yet, it is unclear how well these models perform on biomedical relation extraction (RE). To address this knowledge gap, we explore patterns in the performance of OpenAI LLMs across a diverse sampling of RE tasks.
Methods: We use OpenAI GPT-4-turbo and their reasoning model o1 to conduct end-to-end RE experiments on seven datasets. We use the JSON generation capabilities of GPT models to generate structured output in two ways: (1) by defining an explicit schema describing the structure of relations, and (2) using a setting that infers the structure from the prompt language.
Results: Our work is the first to study and compare the performance of the GPT-4 and o1 for the end-to-end zero-shot biomedical RE task across a broad array of datasets. We found the zero-shot performances to be proximal to that of fine-tuned methods. The limitations of this approach are that it performs poorly on instances containing many relations and errs on the boundaries of textual mentions.
Conclusion: Recent large language models exhibit promising zero-shot capabilities in complex biomedical RE tasks, offering competitive performance with reduced dataset curation and NLP modeling needs at the cost of increased computing, potentially increasing medical community accessibility. Addressing the limitations we identify could further boost reliability. The code, data, and prompts for all our experiments are publicly available: https://github.com/bionlproc/ZeroShotRE
@@ 1 INTRODUCTION
Do we need training data to perform relation extraction (RE)? Since ChatGPT was introduced in December 2022, this has been a prominent question on minds of many NLP researchers, especially that focus on structured information extraction from biomedical or clinical text. With the recent successes of zero-shot research in other areas of NLP, the complex task of relation extraction is ripe for investigation, and biomedicine is a particularly compelling domain.
RE is the high-value NLP task of identifying semantic relationships between entities within text. Consider the following sentence taken from the drug combination extraction (DCE) [1] dataset: “Furthermore, in non-metastatic castration-resistant prostate cancer (M0 CRPC), two second-generation anti-androgens, apalutamide, and enzalutamide, when used in combination with ADT, have demonstrated a significant benefit in metastasis-free survival.” Two beneficial drug combinations are described here: (1) apalutamide with ADT, and (2) enzalutamide with ADT. In RE, we want to extract these relations into a structured form; in this case a tuple of drugs administered in combination along with a signifier of the normative effect of the drug:
@@
Thus, RE can be viewed as the conversion of unstructured data into structured data representing relationships between entities. In biomedicine, entities of interest are mainly genes, mutations, proteins, chemicals, drugs, diseases, and symptoms. The relationships that can hold between them are myriad, but some obvious relationships of importance are drug interactions, protein interactions, disease-causing mutations and chemicals, drug side effects, and disease-treating drugs. Typical texts used for mining biomedical relations are journal abstracts and clinical notes.
@@
Biomedical publications and clinical notes are being generated at breakneck speed — PubMed indexes over 35 million articles and over three thousand more are indexed daily. Biomedical relations are so valuable that there are teams of workers employed to read biomedical text and populate databases with them. With so much text available, automating RE would allow us to mine biomedical relations at scale, rapidly enlarging databases.
@@
Traditionally, RE has been conducted by fine-tuning with hundreds to thousands of examples. In this paradigm, every narrow RE task requires a dataset to be curated for it. Dataset curation for biomedical RE is a laborious process—annotators need highly specialized knowledge, it takes time to develop clear and consistent annotation guidelines, and annotators need to be trained on the guidelines [2, 3]. The process is so laborious for annotators that, typically, named entity recognition and RE tools are used to make suggestions to annotators to speed up the process [2, 3]. For all these reasons, few-shot (FS) and zero-shot (ZS) RE are desirable. They remove the need for a training dataset, which is usually the largest in a training/validation/test split. If modeling choices do not need to be made and if performance need not be measured, validation and test sets could be omitted as well. The promise of high quality low-resource RE is a proliferation of databases and an increase in their richness and reliability.
@@
Generative language models (LMs), especially large ones termed LLMs, have grown to dominate NLP research activity in recent years, with new multi-billion parameter models being released regularly. Owing to large amounts of diverse training data, vast quantities of parameters, and techniques that align models with human goals like instruction finetuning [4, 5, 6] and RLHF [7], these models can now perform a wide array of tasks, in a ZS manner, with impressive results. However, LLMs are not adept at producing long output in a consistent format, a key challenge when converting generated text into structured relations, without specific guiding mechanisms. To address this, previous studies have employed two main strategies: (1) prompting the language model to generate text in a predetermined format, followed by the use of predefined regular expressions to extract relation components [8, 9], or (2) directly specifying a structured output within the prompt itself [10]. These approaches have shown promise, particularly when the model undergoes finetuning [8, 9, 10]. However, Wadhwa et al.  [10] found that even for few-shot sentence-level RE, GPT-3.5 often generates plausible relations that, while recognizable to humans as correct, do not precisely match the gold standard relations. This discrepancy should be expected to be more pronounced in document-level ZS settings. Without fine-tuning, fulfilling synergistically demanding requirements of long, exact, consistent extraction from text becomes significantly more challenging.
@@
Zero-shot relation extraction (ZSRE) poses a unique challenge to LLMs because it requires the generation of long, exact text in a consistent format. This makes it much more challenging to generate an exact output than most other exact-output tasks NLP researchers have been tackling, such as question answering (QA), where the answers are a single token or a phrase [11, 12, 13]. Also, ZS generation has been successful at tackling problems where long text must be generated, like summarizing [11, 12, 13]. In such tasks, there is no single correct generation, so the fact that LLMs produce diverse output is not a problem; these are often evaluated by humans or judgment by other more LLMs (though this practice is controversial). But in the real-world scenario of RE from a document, there may be many relations present (the BioRED [2] dataset contains abstracts with $>50$ relations), and the LM must generate long text with stringent requirements on what the text must consist of.
@@
Because of the aforementioned challenges, LLMs generate relations that are correct to a human but do not match annotations in the test dataset exactly, which artificially deflates calculated performance. For example, if an annotated relation contains the entity hypertensive as the disease in a relation and an LM extracts hypertension, this would be considered incorrect in the usual performance evaluation. Wadhwa et al.[10] deal with this problem using manual evaluation of their relation extraction systems, but this is not desirable because (1) it is expensive and (2) because we want a system that can be used at scale to populate databases automatically, without human intervention.
Given the effectiveness of ZS generation in other NLP tasks, in this paper, we investigate its utility in the high-value task of biomedical RE. We comprehensively test the effectiveness of the OpenAI GPT-4 [13] and OpenAI o1 [14] on seven RE datasets that vary in domain, length of text, diversity of entity and relation types, whether relations are entity-level (EL) or mention-level (ML), and whether relations are described across multiple sentences. We use new features from the OpenAI API that yield structured output. We conduct a thorough error analysis to determine the strengths and weaknesses of this approach. The code, datasets, and the LLM prompts for all our experiments are publicly available for benchmarking ZS end-to-end biomedical RE: https://github.com/bionlproc/ZeroShotRE
2 RELATED WORK
@@
Most relation classification (RC) and RE***We distinguish RC, where gold entities/mentions are available, from RE, where entities must be extracted methods traditionally focus on constructing embeddings for candidate relations, followed by a classification process. A parallel line of research has developed around the use of copy-mechanisms in a sequence-to-sequence (seq2seq) framework. Seq2seq tasks involve generating an output/target sequence from a given input/source sequence. This methodology is predominantly favored in areas like machine translation, where the format aligns naturally with the task. However, researchers have adapted RE to fit into the seq2seq paradigm by transforming structured relations into predefined sequences of tokens [15, 16, 17, 18, 19]. For instance, Giorgi et al. [19] transform the relation {gene: ESR1, disease: schizophrenia, predicate:association} into the sequence ESR1 @gene schizophrenia @disease @associationThis is a slight adaptation from the original paper, simplified for clarity. Subsequently, models are trained to generate such sequences, and decoding them becomes straightforward. The key to the success of these models across various architectures lies in the incorporation of copy mechanisms. In the context of copy mechanism-based RE, the fundamental component is an LSTM that, at each time step, opts to select either a token from the source sequence or a limited additional vocabulary, such as punctuation or special tokens like @gene.
@@
In more recent developments, the use of LLMs has emerged as a novel seq2seq approach for RE [8, 9]. The authors of BioGPT, for example, have fine-tuned their model using soft prompts to generate relations within natural language sentences, such as The relation between <head entity> and <tail entity> is <relation type>. These constructs with place holders for entities and relation types (also called predicates) are often called output templates. The filled-in output template is then processed using regular expressions to extract the relations from the LM’s generations. This method presents a significant advantage over traditional relation representation and copy-mechanism approaches primarily because it does not require mention annotations during training. Such a feature reduces the workload for annotators on additional datasets, as they can shift their focus solely to relation annotation rather than annotating every entity mention. Building on this, Wadhwa et al. [10] modified this approach by designing target sequences as Python-interpretable tuples of relations, rather than in the form of natural sentences, for sentence-level RE tasks.
@@
The remarkable performance of large, human-aligned language models in FS and ZS tasks has sparked interest in exploring their potential for low-resource RE. This emerging area of research particularly focuses on the capabilities of OpenAI’s GPT models. Wadhwa et al. [10] investigate the use of the instruction-finetuned GPT-3.5 for sentence-level biomedical RE. Their FS in-context learning experiments yield results that are competitive with state-of-the-art approaches. In a similar vein, Wang et al. [20] applied GPT-3.5 for sentence-level RC. Further advancing this line of inquiry, Jahan et al. [21] conduct RE experiments using both GPT-3.5 and GPT-4, testing them on two RE dataset test sets, though in one they filter out all examples with no relations.
@@
Our research distinguishes itself from prior studies in several ways. Previous research predominantly explored general biomedical tasks — a valuable effort — but restricting the study to one or two datasets is insufficient to explore the intricacies of RE. Our experiments encompass a broader spectrum of biomedical RE tasks, across seven datasets. We employ datasets that include relations confined within single sentences as well as those with relations spanning across multiple sentences. Moreover, some datasets we study feature relations between entities, while others contain relations between mentions of entities. This variety introduces a range of complexities, including varying levels of difficulty and differing quantities of relations within each text. Such diversity underlines that, although these tasks are all categorized under RE, they each pose unique challenges to RE methodologies. Another significant aspect of our work is the comprehensive evaluation of performance across entire test sets, facilitating direct comparisons with other studies, whether they are fine-tuned, FS, or ZS. Next, we use features of the OpenAI API that restrict model output to be a valid JSON object, facilitating direct extraction of structured output. Lastly, our study undertakes an in-depth analysis of the specific challenges posed by zero-shot generative RE.
@@ 3.1 Task Definition
Let $\left(x,y\right)$ be an example in the test dataset, where $x$ is text and $y$ is the set of annotated relations expressed within $x$. Depending on the dataset, the text and relations may have different structures. $x$ is most commonly a title-abstract pair but maybe a sentence along with a larger passage containing it or a single string of text. $y$ depends on the dataset as well. In general, a relation consists of a set of typed entities along with a relation type connecting them. In most datasets, relations hold between two entities, but in the DCE [1] dataset, they hold between a variable number of entities. Entities and relations are typed, but the number of types varies by dataset.
@@ 3.2 Extraction
Most previous work on generative RE has relied on traditional supervision. In such scenarios, the choice of template is of moderate importance, because $\mathcal{M}$ is finetuned to learn the nature of the problem and the structure of the output. Without a supervision signal, it is challenging to guide $\mathcal{M}$ to (1) understand the nature of the problem and (2) output relations in a consistently structured form. The latter is important for biomedicine if RE is to be automated, and important for research as it permits performance metrics to be calculated. To address these challenges, $T$ adds a complete description of the RE task as well as instructions to produce output in the form of a JSON object, a description of the format the JSON object should take, and an example of what a filled-in JSON could look like.
Given that producing consistent structured output from a language model that has been principally trained to produce natural language is a problem faced in information extraction in general, a few attempts have been made to solve it [22, 23] or produce a structured output posthoc [24]. Given that we use GPT-4[13] and o1 [14] as $\mathcal{M}$ and the fact that OpenAI recently added functionality in their API for obtaining JSON objects as output in two different modes, we use their tools. The first tool they developed requires the user to provide a schema (in the form of a JSON object) delineating the structure the output JSON should exhibit. The more recent tool infers the schema from the prompt. We refer to these modes as explicit and inferred modes and experiment with both on GPT-4. For OpenAI o1 [14], we test only the mode that performed better on GPT-4, due to budgetary constraints. Unlike GPT-4, the o1 model is designed with test-time chain-of-thought based reasoning ability; it learns to “recognize and correct its mistakes” and “ break down tricky steps into simpler ones."https://openai.com/index/learning-to-reason-with-llms/ The overall flow of our work is shown in Figure 1.
Figure 1: Overall flow of the zero-shot relation extraction using OpenAI API.

### v3

@@ Abstract
Extracting relations from scientific literature is a fundamental task in biomedical NLP because entities and relations among them drive hypothesis generation and knowledge discovery. As literature grows rapidly, relation extraction (RE) is indispensable to curate knowledge graphs to be used as computable structured and symbolic representations. With the rise of LLMs, it is pertinent to examine if it is better to skip tailoring supervised RE methods, save annotation burden, and just use zero shot RE (ZSRE) via LLM API calls. In this paper, we propose a benchmark with seven biomedical RE datasets with interesting characteristics and evaluate three Open AI models (GPT-4, o1, and GPT-OSS-120B) for end-to-end ZSRE. We show that LLM-based ZSRE is inching closer to supervised methods in performances on some datasets but still struggles on complex inputs expressing multiple relations with different predicates. Our error analysis reveals scope for improvements.
A benchmark for end-to-end zero-shot biomedical relation extraction
with LLMs: experiments with OpenAI models
@@ 1 Introduction
Do we need training data to perform relation extraction (RE)? Since ChatGPT was introduced in December 2022, this has been a prominent question on minds of many NLP researchers, especially those that focus on structured information extraction from biomedical literature. With the recent success of zero-shot methods in other areas of NLP, RE is ripe for investigation, and biomedicine is a particularly compelling domain as relations are central to knowledge discovery.
RE is the high-value NLP task of identifying semantic relationships between entities within text. Consider the following sentence taken from the drug combination extraction (DCE) Tiktinsky et al. (2022) dataset: “Furthermore, in non-metastatic castration-resistant prostate cancer (M0 CRPC), two second-generation anti-androgens, apalutamide, and enzalutamide, when used in combination with ADT, have demonstrated a significant benefit in metastasis-free survival.” Two beneficial drug combinations are described here: (1) apalutamide with ADT, and (2) enzalutamide with ADT. In RE, we want to extract these relations into a structured form; in this case a tuple of drugs administered in combination along with a signifier of the normative effect of the drug:
@@
Thus, RE can be viewed as the conversion of unstructured data into structured data representing relationships between entities. In biomedicine, entities of interest are mainly genes, mutations, proteins, chemicals, drugs, diseases, and symptoms. The relationships that can hold between them are myriad, but some obvious relationships of importance are drug interactions, protein interactions, disease-causing mutations and chemicals, drug side effects, and disease-treating drugs. Many RE efforts assume that the entities and their types are already provided as part of the input. They do RE by giving the input text and a pair of entity spans in it and ask for the relation type linking them. The technical name for this would be relation classification (RC). However, for RE to be fully automated and evaluated fairly, the input must be just free text and the burden of (a) spotting the entities and (b) linking them with predicates, both fall on the method. This is called end-to-end RE (E2E RE) and is much harder than RE when entities are pre-annotated. For the rest of this paper, whenever we refer to RE we mean the end-to-end variety.
@@
Biomedical publications are being generated at breakneck speed — PubMed indexes nearly 40 million articles and over four thousand more are indexed daily. Biomedical relations are so valuable that there are teams of workers employed to read biomedical text and populate databases with them. With so much text available, automating RE would allow us to mine biomedical relations at scale, rapidly enlarging databases.
1.1 Zero-shot RE
@@
Traditionally, RE has been conducted by fine-tuning with hundreds to thousands of examples. In this paradigm, every narrow RE task requires a dataset to be curated for it. Dataset curation for biomedical RE is a laborious process — annotators need highly specialized knowledge, it takes time to develop clear and consistent annotation guidelines, and annotators need to be trained on the guidelines Luo et al. (2022a); Li et al. (2016). The process is so laborious for annotators that, typically, named entity recognition and RE tools are used to make suggestions to annotators to speed up the process Luo et al. (2022a); Li et al. (2016). For all these reasons, few-shot (FS) and zero-shot (ZS) RE are desirable. They remove the need for a training dataset, which is usually the largest in a training/validation/test split. If modeling choices do not need to be made and if performance need not be measured, validation and test sets could be omitted as well. The promise of high quality low-resource RE is a proliferation of databases and an increase in their richness and reliability.
@@
Generative large language models (LLMs) have grown to dominate NLP research activity in recent years, with new multi-billion parameter models being released regularly. Owing to large amounts of diverse training data, vast quantities of parameters, and techniques that align models with human goals like instruction finetuning Ouyang et al. (2022a); Wei et al. (2022); Sanh et al. and RLHF Ouyang et al. (2022b), these models can now perform a wide array of tasks, in a ZS manner, with impressive results. However, LLMs are not adept at producing long output in a consistent format, a key challenge when converting generated text into structured relations, without specific guiding mechanisms. To address this, previous studies have employed two main strategies: (1) prompting the language model to generate text in a predetermined format, followed by the use of predefined regular expressions to extract relation components Luo et al. (2022b); Gupta et al. (2025), or (2) directly specifying a structured output within the prompt itself Wadhwa et al. (2023). Though these approaches have shown promise, Wadhwa et al. (2023) found that even for few-shot sentence-level RE, GPT-3.5 often generates plausible relations that, while recognizable to humans as correct, do not precisely match the gold standard relations. This discrepancy should be expected to be more pronounced in document-level ZS settings. Without fine-tuning, fulfilling synergistically demanding requirements of long, exact, consistent extraction from text becomes more challenging.
@@
Zero-shot relation extraction (ZSRE) poses a unique challenge to LLMs because it requires the generation of long, exact text in a consistent format. This makes it much more challenging to generate an exact output than most other exact-output tasks NLP researchers have been tackling, such as question answering (QA), where the answers are a single token or a phrase Touvron et al. (2023); Wang et al. (2023a); Achiam et al. (2023). Also, ZS generation has been successful at tackling problems where long text must be generated, like summarizing Touvron et al. (2023); Wang et al. (2023a); Achiam et al. (2023). In such tasks, there is no single correct generation, so the fact that LLMs produce diverse output is not a problem; these are often evaluated by humans or judgment by other more LLMs (though this practice is controversial). But in the real-world scenario of RE from a document, there may be many relations present (the BioRED Luo et al. (2022a) dataset contains abstracts with >50>50 relations), and the LM must generate long text with stringent requirements on what the text must consist of.
@@
Due to the aforementioned challenges, LLMs generate relations that are correct to a human but do not match annotations in the test dataset exactly, which artificially deflates calculated performance. For example, if an annotated relation contains the entity hypertensive as the disease in a relation and an LM extracts hypertension, this would be considered incorrect in the usual performance evaluation. Wadhwa et al. (2023) deal with this problem using manual evaluation of their relation extraction systems, but this is not desirable because (1) it is expensive and (2) because we want a system that can be used at scale to populate databases automatically, without human intervention.
1.2 Related work
@@
Most RE methods focus on constructing embeddings for candidate relations, followed by a classification step. A parallel line of research has developed around the use of copy-mechanisms in a sequence-to-sequence (seq2seq) framework. Seq2seq tasks involve generating an output/target sequence from a given input/source sequence. This method is predominantly favored in areas like machine translation, where the format aligns naturally with the task. However, researchers have adapted RE to fit into the seq2seq paradigm by transforming structured relations into predefined sequences of tokens Zeng et al. (2018); Zhang et al. (2020); Nayak and Ng (2020); Zeng et al. (2020); Giorgi et al. (2022). For instance, Giorgi et al. (2022) transform the relation {gene: ESR1, disease: schizophrenia, predicate:association} into the sequence ESR1 @gene schizophrenia @disease @association***This is a slight adaptation from the original paper, simplified for clarity. Subsequently, models are trained to generate such sequences, and decoding them becomes straightforward. The key to the success of these models across various architectures lies in the incorporation of copy mechanisms. In the context of copy mechanism-based RE, the fundamental component is an LSTM that, at each time step, opts to select either a token from the source sequence or a limited additional vocabulary, such as punctuation or special tokens like @gene.
@@
In more recent developments, the use of LLMs has emerged as a novel seq2seq approach for RE Luo et al. (2022b); Gupta et al. (2025). The authors of BioGPT, for example, have fine-tuned their model using soft prompts to generate relations within natural language sentences, such as The relation between <head entity> and <tail entity> is <relation type>. These constructs with place holders for entities and relation types (also called predicates) are often called output templates. The filled-in output template is then processed using regular expressions to extract the relations from the LM’s generations. This method presents a significant advantage over traditional relation representation and copy-mechanism approaches primarily because it does not require mention annotations during training. Such a feature reduces the workload for annotators on additional datasets, as they can shift their focus solely to relation annotation rather than annotating every entity mention. Building on this, Wadhwa et al. (2023) modified this approach by designing target sequences as Python-interpretable tuples of relations, rather than in the form of natural sentences, for sentence-level RE tasks.
@@
The remarkable performance of large, human-aligned language models in FS and ZS tasks has sparked interest in exploring their potential for low-resource RE. This emerging area of research particularly focuses on the capabilities of OpenAI’s GPT models. Wadhwa et al. (2023) investigate the use of the instruction-finetuned GPT-3.5 for sentence-level biomedical RE. Their FS in-context learning experiments yield results that are competitive with state-of-the-art approaches. In a similar vein, Wang et al. (2023b) applied GPT-3.5 for sentence-level RC. Further advancing this line of inquiry, Jahan et al. (2023) conduct RE experiments using both GPT-3.5 and GPT-4, testing them on two RE dataset test sets, though in one they filter out all examples with no relations.
1.3 Our contributions
Given the effectiveness of ZS generation in other NLP tasks, in this paper, we investigate its utility in the high-value task of biomedical RE. We comprehensively test the effectiveness of OpenAI GPT-4 Achiam et al. (2023), OpenAI o1 Jaech et al. (2024), and the open-weights GPT-OSS OpenAI (2025) on seven RE datasets that vary in domain, length of text, diversity of entity and relation types, whether relations are entity-level (EL) or mention-level (ML), and whether relations are described across multiple sentences. We analyze model errors to determine the strengths and weaknesses of this approach. The code, datasets, and the LLM prompts for all our experiments are available here: https://github.com/bionlproc/ZeroShotRE.
@@ 1.3 Our contributions
Our research differs from prior studies in several ways. Previous research predominantly explored general biomedical tasks — a valuable effort — but restricting the study to one or two datasets is insufficient to explore the intricacies of RE. Our experiments encompass a broader spectrum of biomedical RE tasks, across seven datasets. We employ datasets that include relations confined within single sentences as well as those with relations spanning across multiple sentences. Moreover, some datasets we study feature relations between entities, while others contain relations between mentions of entities. This variety introduces a range of complexities, including varying levels of difficulty and differing quantities of relations within each text. Such diversity underlines that, although these tasks are all categorized under RE, they each pose unique challenges to RE methodologies. Another crucial aspect of our work is the comprehensive evaluation of performance across entire test sets, facilitating direct comparisons with other studies, whether they are fine-tuned, FS, or ZS.
@@ 2.1 Task definition
Let $\left(x,y\right)$ be an example in the test dataset, where $x$ is text and $y$ is the set of annotated relations expressed within $x$. Depending on the dataset, the text and relations may have different structures. $x$ is most commonly a title-abstract pair but maybe a sentence along with a larger passage containing it or a single string of text. $y$ depends on the dataset as well. In general, a relation consists of a set of typed entities along with a relation type connecting them. In most datasets, relations hold between two entities, but in the DCE Tiktinsky et al. (2022) dataset, they hold between a variable number of entities. Entities and relations are typed, but the number of types varies by dataset.
@@ 2.2 Extraction
Most prior work on generative RE has relied on traditional supervision. In such scenarios, the choice of template is of moderate importance, because $\mathcal{M}$ is finetuned to learn the nature of the problem and the structure of the output. Without a supervision signal, it is challenging to guide $\mathcal{M}$ to (1) understand the nature of the problem and (2) output relations in a consistently structured form. The latter is important for biomedicine if RE is to be automated, and important for research as it permits performance metrics to be calculated. To address these challenges, $T$ adds a complete description of the RE task as well as instructions to produce output in the form of a JSON object, a description of the format the JSON object should take, and an example of what a filled-in JSON could look like.
Given that producing consistent structured output from a language model that has been principally trained to produce natural language is a problem faced in information extraction in general, a few attempts have been made to solve it Newhouse (2023); Sengottuvelu (2023) or produce a structured output posthoc Yurtsev (2023). Given that we use GPT-4 Achiam et al. (2023), o1 Jaech et al. (2024) and GPT-OSS OpenAI (2025) as $\mathcal{M}$ and the fact that OpenAI added functionality in their API for obtaining JSON objects as output in two different modes, we use their tools. The first tool they developed requires the user to provide a schema (in the form of a JSON object) delineating the structure the output JSON should exhibit. The more recent tool infers the schema from the prompt. We refer to these modes as explicit and inferred modes and experiment with both on GPT-4. For OpenAI o1 and GPT-OSS Jaech et al. (2024); OpenAI (2025), we test only the mode that performed better on GPT-4, due to budgetary constraints. Unlike GPT-4, the other two models are designed with test-time chain-of-thought based reasoning ability; it learns to “recognize and correct its mistakes” and “ break down tricky steps into simpler ones."https://openai.com/index/learning-to-reason-with-llms/